a commercial jet and a private airplane leave the same airfield at the same time and travel in opposite directions at average speeds of 880 km/hr and 450 km/hr respectively. In how many minutes will they be 1463 km apart?
Answer by mananth(16949)

Commercial jet -------- 880 km/h
Private plane -------- 450 km/h

they are traveling away from each other
so add up speeds
450 + 880 = 1330 km/h
Distance = 1463 km
time = d/r
time= 1463 / 1330
Time = 1.1 h

1.1 * 60 = 66 minutes
m.ananth@hotmail.ca